OpenStudy (anonymous):

50,000/65,000 reduce to lowest terms and how do u get answer?>

OpenStudy (mathteacher1729):

First, anytime you have stuff with lots of zeros at the end, you can just chop off the zeros: 100/100 = 1 10/1,000 = 1/100 1,000,000/1,000 = 1,000/1 etc. Now, let's say we have 20,000/85,000 .... this reduces to 20/85. Now just reduce 20/85 by factoring and canceling like factors. :) From this, can you see where to go with your prob?

OpenStudy (anonymous):

for some reason these are soo confusing lets see.. like factors would be....5?

OpenStudy (mathteacher1729):

"like factors would be 5" -- yes. :) Any number which ends in 0 or 5 is automatically divisible by 5.

OpenStudy (anonymous):

cancel out zeros: 50/65 divide by 5: 10/13

OpenStudy (anonymous):

so the answer is yaya i was right ok for some reason i did the caculator thing online and got different thats scary..

OpenStudy (mathteacher1729):

roseanne, I left a message on one of your earlier posts about http://www.mathtv.com -- you might find that website very helpful. (click basic math > and then there's tons of stuff about ratio, percents, fractions, etc.)

OpenStudy (anonymous):

oh thank you!!!